OpenStudy (anonymous):

The lab table has a perimeter of 214 inches. The length is 13 inches less than 3 times the width. How many inches long is the table?

OpenStudy (johnweldon1993):

Well you know the perimeter would be 2L + 2W (rectangle) We know the length (L) is 13 less than 3 times the width (W) so that means... L = 3W - 13 Since we know that...we can substitute that into our equation for L so 2L + 2W = 214 would become 2(3W - 13) + 2W = 214 Can you solve this from here?

OpenStudy (anonymous):

x = 30in, the width. The length is 77in 3%2A30-13 and... 60in+%2B+154in+=214in

OpenStudy (anonymous):

L=77?

OpenStudy (johnweldon1993):

That is correct :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

plz